OpenStudy (usukidoll):

Suppose the function \(f \rightarrow X \rightarrow Y\) is onto. Prove or disprove that the induced map \(\bar f^{-1} \mathcal P \left({Y}\right) \rightarrow \mathcal P \left({X}\right)\) is onto. As a first step, make sure to state what it means that a function is onto. The function \(f: X \rightarrow Y\) is onto which means it's a surjection, and the definition is \((\forall Y) (\exists X) (f(x) = y)\)

OpenStudy (usukidoll):

@wio do I just provide an example that would prove that this is a surjection?

OpenStudy (usukidoll):

like... what if we let X = [a,b,c,d] and Y = [1.2,3] then we could find the pairings to confirm that \[\bar f^{-1} \] is indeed a function from powerset (y) to powerset (x)

OpenStudy (usukidoll):

f(a) = 1 f(b) = 3 f(c) = 2 f(d) = 3

OpenStudy (anonymous):

induced map?

OpenStudy (usukidoll):

yes

OpenStudy (usukidoll):

isn't that the union of all elements?

OpenStudy (anonymous):

I need a clear definition of induced map

OpenStudy (usukidoll):

k let me check in my book

OpenStudy (usukidoll):

Let \(f: X \rightarrow Y\). The set function induced by \(f\) is the function \(\bar f \mathcal P \left({X}\right) \rightarrow \mathcal P \left({Y}\right) \) defined by the rule that for all \(A \in \mathcal P \left({X}\right) \). \(\bar f(A) =[y \in Y : (\exists x \in A) [f(x) =y]]=[f(x): x \in A]\)

OpenStudy (usukidoll):

hmm might work if the bar wasn't an inverse.. there's another definition with f bar inverse

OpenStudy (usukidoll):

def... 5.3.8 Let \( f : X \rightarrow Y\). For each set \(B \in \mathcal P \left({Y}\right) . \), define the function \(\bar f^{-1} : \mathcal P \left({Y}\right) \rightarrow \mathcal P \left({X}\right)\) by \(\bar f^{-1} (B) =[x \in X: f(x) \in B]\) . .

OpenStudy (anonymous):

What about \(\bar{f} ^{ -1}(\emptyset)\)?

OpenStudy (usukidoll):

wouldn't that be an emptyset?! ^

OpenStudy (usukidoll):

prove or disprove...

OpenStudy (anonymous):

Okay how about this....

OpenStudy (usukidoll):

so.. if we pick wrong..oh boy

OpenStudy (anonymous):

We should try to find a counter example.

OpenStudy (anonymous):

I do have one idea though...

OpenStudy (usukidoll):

ok... so if isn't not a surjection then f(x) can't equal to y

OpenStudy (usukidoll):

*if it

OpenStudy (anonymous):

\[ \bar f^{-1}(\{1,2\}) = \{1,2,3\} \]and\[ \bar f^{-1}(\{2\}) = \{3\} \]

OpenStudy (anonymous):

I think from your counter example, there is no way to get \(\bar f^{-1}(S) = \{2,3\}\) since if you want \(2\) then you're going to get \(1\) as well.

OpenStudy (anonymous):

So while \(\bar f^{-1}\) is a function, it isn't onto.

OpenStudy (usukidoll):

ok but shouldn't we have the f first and then the f inverse second.. and finally the f bar inverse.. because my book has it as f: [1,2,3] -> [a,b,c] and it was defined by f(1) = a f(2) = a and f(3) = b ... so the inverses are f^-1(a) = 1 f^-1(a) = 2 and f^-1(b) = 3

OpenStudy (usukidoll):

oops that's f inverse isn't a function D:

OpenStudy (usukidoll):

what if we want to prove it.. what would we need to do? just wondering

OpenStudy (anonymous):

X={1,2,3} Y={a,b} f(1) = a f(2) = a f(3) = b fbar({1,2,3})={a,b} fbar({2,3})={a,b} fbar^-1 {a,b}={1,2,3} as wio pointed out, fbar^-1 will never equal {2,3}, an element of p(X) thus fbar^-1 is not onto

OpenStudy (usukidoll):

oh I see it... they don't equal at all...so it's not a surjection ..

OpenStudy (usukidoll):

would anyone like to check my induction proof? I'm kind of whizzed that I got it wrong because I swear I've done everything right.

OpenStudy (anonymous):

\[ \bar f^{-1}(\emptyset) = \emptyset \\ \bar f^{-1}(\{a\}) = \{1,2\} \\ \bar f^{-1}(\{b\}) = \{3\} \\ \bar f^{-1}(\{a,b\}) = \{1,2,3\} \]You can't find \(\{1\}\), \(\{2\}\), \(\{1,3\}\) or \(\{2,3\}\)

OpenStudy (anonymous):

Also, when you think about it, if \( \|\mathcal P(Y)\| < \|\mathcal P(X)\| \) then it is impossible for it to be onto.

OpenStudy (usukidoll):

k... new question..